OpenStudy (anonymous):

How can I write a quadratic function in vertex form? vertex@(-1,-6) graph contains the point (0,-8) vertex@(5,3) graph contains the points (4,6) vertex@(-3,-6) graph contains the point (3,11) x-inter. of graph is 8 and -4; y-inter. is -12

OpenStudy (anonymous):

The general vertex form is: y = a(x - h)^2 + k where (h, k) is your vertex, so looking at the first one: y = a(x - (-1))^2 + (-6) = a(x + 1) - 6 Substitute the other point to get the value for "a" -8 = a(0 + 1) - 6 -> -8 = a - 6 -> a = -2 Substitute "a" back in along with h and k y = -2(x + 1)^2 -6

OpenStudy (anonymous):

Okay, I think i got it.

OpenStudy (anonymous):

Good news!